# ForestSplats: Deformable transient field for Gaussian Splatting in the Wild

### v2

@@ 6 Evaluation
As shown in Fig. 4 and Tab. 2, our method shows favorable performance compared to prior methods. In Fig. 4, we observe that CR-NeRF [40], 3D-GS [7], and GS-W [44] frequently fail to capture fine details in scenarios with appearance variations and struggle to handle transient elements. Unlike existing methods, WildGaussians [10] and Wild-GS [37] leverage semantic features from pre-trained VFM [39, 20] to handle the transient elements. In contrast, our ForestSplats effectively decomposes transient elements from 2D scenes by leveraging transient field and a superpixel-aware mask without VFM, demonstrating competitive performance, as shown in Tab. 2. More details are in the supplements.
@@
Comparison on the NeRF On-the-go Dataset. Furthermore, we also show the results for the NeRF On-the-go dataset [22]. Our method also achieves competitive performance against existing SOTA methods across various occlusion levels while enabling real-time rendering without VFM in Tab. 1. Compared to RobustNeRF [23], and NeRF On-the-go [22], our method exhibits faster training and rendering speeds while mitigating negative occluders. RobustNeRF [23] considers photometric errors and the small size of the sampled patches to overlook complex structures and high occlusion scenarios. Specifically, WildGaussians [10] and SpotLessSplats [24] effectively remove transient elements by considering semantic features. Although leveraging semantic features facilitates handling transient elements, distinguishing semantic features similar to static elements remains challenging. However, our method explicitly decomposes static and transient elements across different levels of occluder scenarios, as shown in Fig. 5. Furthermore, compared to the prior methods without VFM, our method outperforms average performance in terms of PSNR and SSIM. We provide more results in the supplements.
7 Ablation Study
Efficiency of Deformable Transient Field.
As shown in Tab. 3, the deformable transient field significantly reduces the CPU memory usage (MB). Specifically in the Sacre Coeur, although transient embedding per image is considered, the deformable transient field achieves over an 80% improvement in CPU memory efficiency (MB). Compared to existing methods [33, 14], the deformable transient field enables all transient Gaussians to be loaded onto the GPU, avoiding unnecessary context switching between the CPU and GPU. Furthermore, we also analyze the performance based on the number of transient Gaussians as shown in Fig. 7-(a). We observe that the number of transient Gaussians primarily aims to occupy the regions with transient elements, which results in a slight effect on the representation of the static field. Moreover, we compare the transient field with and without the deformation MLP, as shown in Tab. 4. These results indicate that the deformable transient field not only significantly reduces memory usage (MB) but also facilitates scalability, particularly in large-scale scenarios.
@@ 7 Ablation Study
Effectiveness of superpixel-aware mask.
We conducted an ablation study to assess the effectiveness of a superpixel-aware mask in handling transient elements. As shown in Fig. 8, our method explicitly decomposes distractors from the static scene and represents transient objects in various scenarios, handling varying levels of occlusion. Furthermore, we analyze the effect of the number of superpixels in Fig. 9. According to the results, a small number of superpixels has minimal effect on the transient mask. Meanwhile, a large number of superpixels may lead to excessive masking of irrelevant areas. Thus, we set the number of segments to 20 as an optimal number. Moreover, we compare the quality of the transient mask in Fig. 10-(a). Existing methods [40, 44] rely on photometric errors. In contrast, superpixel-aware masks leverage superpixels, which consider both color and spatial information, to enhance mask quality without additional memory cost. Furthermore, the superpixel-aware mask serves as a plug-and-play module that combines with existing transient mask methods, showing improved performance, as shown in Fig. 10-(b). The results indicate that the superpixel-aware mask improves performance, achieving an average of 1.47 dB in PSNR and 0.022 in SSIM. Furthermore, we demonstrate a multi-stage training scheme that consistently improves performance in Fig. 7-(b). More results are provided in the supplements.
Discussion on uncertainty-aware densification.
We use uncertainty-aware densification (UAD) only in the static field to avoid the generation of static Gaussians at the boundaries of distractors during optimization. To analyze this, we investigate an ablation study with and without UAD to see its effect on novel-view synthesis. As shown in Fig. 11, we observe that the centroids of static Gaussians are primarily positioned on static elements, mitigating the artifacts in novel view synthesis.
Analysis of each component.
For further analysis, we evaluate each component of our method to demonstrate the effectiveness: Deformable transient field (DTF), Superpixel-aware mask (SAM), Multi-stage training scheme (MST), and Uncertainty-aware densification (UAD) as shown in Tab. 5. The results indicate that a superpixel-aware mask, which considers the photometric errors and the superpixels, combined with a multi-stage training scheme, consistently improves performance by effectively removing transient elements in scenes. Furthermore, we show qualitative results on the Brandenburg Gate and Corner scenes in Fig. 12.
@@ 8 Limitation and Conclusion
Although the transient field does not exhibit high quality due to deformable Gaussians, as shown in Fig. 8, the deformable transient field effectively captures transient elements while showing significant memory efficiency. Another limitation of our forestsplats is its difficulty in generalizing the number of superpixels due to the varying levels of occlusion in datasets, as depicted in Fig. 9.
Conclusion.
@@
In this paper, we introduce ForestSplats, a novel framework that leverages a deformable transient field and a superpixel-aware mask to effectively remove distractors from static scenes. Furthermore, we propose a simple yet effective uncertainty-aware densification to enhance rendering quality and avoid generating static Gaussians on the boundaries of distractors during optimization. Extensive experiments on several datasets demonstrate that our method shows state-of-the-art performance compared to existing methods without a pre-trained model and shows significant memory efficiency to represent transient elements.
Supplementary Material
In this supplementary materials, we provide a detailed description of the more complete experimental settings, along with additional qualitative and quantitative results, including metrics that were not included in the main paper due to the page limitations. Furthermore, we present additional ablation studies, detailed analyses and observations, and provide failure cases. We also discuss potential future directions. The contents are summarized as follows:
  • Sec. 9: More implementation details
  • Sec. 10: More experimental results
  • Sec. 11: More ablation studies
  • Sec. 12: User study
  • Sec. 13: Failure cases and Future works
9 More Implementation details
9.1 Architecture
Our method is based on the WildGaussians [10] codebase. For the deformable transient field, we follow Deform3D-GS [41] for modeling transient elements. We train a deformable MLP $f_{d}$ with the transient field to capture transient elements. Otherwise, we design a color mapping MLP $f_{c}$ consisting of eight fully connected layers with ReLU activations, 256-dimensional hidden layers, and a 256-dimensional output. We use transient embedding per image $t\in\mathbb{R}^{32}$ for deformable MLP $f_{d}$. During inference phase, we only leverage static field for novel-view synthesis.
9.2 Photo Tourism & NeRF On-the-go
For the Photo Tourism dataset [27], we optimize for 200K iterations to train a representation of the entire scene. For the NeRF On-the-go dataset [22], we optimize for 30K iterations. Most hyperparameter settings follow the default codebase [10]. To represent transient elements and handle appearance changes in the Photo Tourism dataset [27], we leverage an appearance MLP, following WildGaussians [10]. Additionally, we introduce a multi-stage training scheme, where $\mathcal{L}_{init}$ is applied up to 80K iterations, followed by $\mathcal{L}_{mid}$ until 120K iterations, and finally, $\mathcal{L}_{total}$ is used for the remaining steps. For the NeRF On-the-go dataset [22], we apply $\mathcal{L}_{init}$ for the first 15K iterations, continue with $L_{mid}$ until 20K iterations, and finally switch to $\mathcal{L}_{total}$.
Scene Method Memory Scene Method Memory
Mountain Desplat [33] 29.03 Fountain Desplat [33] 37.81
HybridGS [14] 44.29 HybridGS [14] 57.68
ForestSplats (Ours) 2.28 ForestSplats (Ours) 2.29
Patio Desplat [33] 22.28 Spot Desplat [33] 41.86
HybridGS [14] 33.99 HybridGS [14] 63.86
ForestSplats (Ours) 2.26 ForestSplats (Ours) 2.27
Table 6: Comparison of memory usage efficiency. We report CPU Memory usage (MB) for the initial number of transient Gaussians.
10 More experimental results
We provided additional experiment results to demonstrate a more comprehensive evaluation of our methods on the Photo Tourism [27] and NeRF On-the-go [22] datasets. Furthermore, we provide additional qualitative results to validate the effectiveness of each component in Fig. 24. Moreover, we highly recommend that the reader watch the several videos on the webpage. Our method achieves consistency and high-quality novel-view synthesis.
Figure 13: Visualization of sample training images. We present occluders in the Photo Tourism and NeRF On-the-go datasets.
10.1 More Photo Tourism results
As illustrated in Fig. 18 and Tab. 9, our methods achieve high-quality rendering and show competitive performance compared to existing methods without VFM. Wild-GS and WildGaussians use semantic features from pre-trained VFM to generalize static scenes. In contrast, our method effectively decomposes transient elements from static scenes without VFM. Furthermore, we gradually interpolate between two appearance embeddings, as shown in Fig. 20. The results indicate the smoothness and consistency of the appearance embedding.
10.2 More NeRF On-the-go results
For NeRF On-the-go dataset [22], we visualize additional results of the remaining scenes due to page limitations as shown in Fig. 19. NeRF On-the-go [22] and WildGaussians [10] tackle transient elements using semantic features from the Vision Foundation model. However, our method addresses transient elements by considering photometric errors and superpixels. The detailed quantitative results demonstrate that ForestSplats outperforms the prior methods in Tab. 10. Specifically, our ForestSplats achieves state-of-the-art performance across five scenes, showing high-quality results.
Scene Method Memory usage (MB) Testing FPS
Static field Transient field
Corner 3D-GS [7] 410.70 0 116
HybridGS [14] 42.20 34.70 160
ForestSplats (Ours) 58.0 2.27 143
Trevi Fountain GS-W [10] 91.21 0 38
ForestSplats (Ours) 30.5 3.39 103
Table 7: Comparison of memory usage (MB) and testing FPS.
Segments PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
10 19.83 0.65 0.324
15 21.47 0.711 0.198
20 22.61 0.781 0.145
25 22.21 0.738 0.199
Table 8: Comparison of the number of superpixels.
11 More ablation studies
11.1 Efficiency of deformable transient field
As shown in Tab. 6 and Tab. 7, we emphasize the advantages of our method, particularly in terms of memory usage and efficiency. Although ForestSplats requires the Deformable MLP and the color mapping MLP, our method efficiently utilizes memory to represent transient elements. Compared to GS-W [44], our method demonstrates remarkable computational efficiency on the Trevi Fountain scene in Tab. 7.
11.2 Effectiveness of superpixel-aware mask
As shown in Fig. 21, superpixel-aware mask effectively captures the transient elements compared to the existing methods. Furthermore, we demonstrate that our method explicitly decomposes the transient elements from the 2D scenes, as provided in Fig. 22. Moreover, we provide additional results for the multi-stage training scheme in Fig. 14. We also provide detailed results for the number of superpixels in Tab. 8. As shown in Fig. 13, the distribution of transient elements varies between the Photo Tourism and NeRF On-the-go datasets. Therefore, adjusting the number of superpixels for each scene could enhance performance.
Figure 14: Comparison of training PSNR across with and without a multi-stage training scheme. MST denotes multi-stage training.
11.3 Proof of positional gradient
We provide a step-by-step proof that the positional gradient is generally directed towards a region of static elements. The gradient is calculated by:
$\mathcal{G}_{i,x}=\frac{\partial\mathcal{L}}{\partial\mu_{i,x}}=\sum_{k=1}^{p}\frac{\partial\mathcal{L}_{k}}{\partial\mu_{i,x}},$ (11)
where $\mathcal{L}$ denotes the rendering loss from $\mathcal{L}_{\mathrm{GS}}$, $p$ is the number of pixels covered by Gaussian $\mathcal{G}_{i}$, and $\mu_{i,x}$ is the pixel-space projection of Gaussian $\mathcal{G}_{i}$ under viewpoint $x$. The computation for static Gaussians is as follows:
$\frac{\partial\mathcal{L}_{k}^{(static)}}{\partial\mu_{i,x}}\propto(1-\mathcal{M}_{S}(x_{k}))\frac{\partial\mathcal{L}_{k}}{\partial\mu_{i,x}}.$ (12)
On the other hand, the positional gradient for transient Gaussians is computed as:
$\frac{\partial\mathcal{L}_{k}^{(transient)}}{\partial\mu_{i,x}}\propto\mathcal{M}_{S}(x_{k})\frac{\partial\mathcal{L}_{k}}{\partial\mu_{i,x}}.$ (13)
Thus, the positional gradient is influenced by the superpixel-aware mask $\mathcal{M}_{S}$. In the Photo Tourism [27] and NeRF On-the-Go datasets [22], transient elements typically constitute less than 30%. Consequently, the gradient tends to be directed toward regions of static elements.
11.4 Discussion of uncertainty-aware densification
We further demonstrate that our uncertainty-aware densification (UAD) effectively removes high-uncertainty static Gaussians, thereby enhancing the consistency and reconstruction quality in Fig. 15 and Fig. 23. Specifically, UAD enhances depth consistency compared to the case without UAD, as illustrated in Fig. 15. Furthermore, we performed an additional evaluation using SIFT-based feature matching on images rendered with and without UAD. As shown in Fig. 23, leveraging UAD ensures that feature correspondences remain consistent across different viewpoints by removing high-uncertainty static Gaussians and avoiding generating static Gaussian in the region of transient elements.
Figure 15: Comparison of depth consistency w/o and w/ UAD.
12 User study
12.1 User evaluation results
We recruited several researchers as survey participants and asked them to complete a questionnaire, as shown in Fig. 25. The questionnaire includes several questions and rendering results comparing our method with baselines on three different scenes. Baseline methods include WildGaussians [10], GS-W [44], NeRF On-the-go [22]. Each participant was asked to evaluate the rendering videos on a scale from 1 (poor) to 5 (excellent). The average scores for each question are visualized in Fig. 17. The results show that our method achieved the highest scores across all aspects, including transient removal, floater reduction, and consistency quality, demonstrating the effectiveness of our approach.
Figure 16: Failure cases of our method. (a) Our method fails to capture fine details in masks within complex scenes. (b) Furthermore, our method fails to generalize to sparse training views.
Figure 17: User Study Results: Average response scores on transient removal, floater reduction, and consistency quality.
13 Failure cases and Future works
13.1 Failure cases
Our method shows impressive results in capturing transient elements. However, similar to prior methods, our method struggles to accurately mask transient elements, as shown in Fig. 16-(a) due to complex real-world environments. Furthermore, as shown in Fig. 16-(b), our method often fails to generalize in novel-view synthesis, especially in scenarios with sparse training views.
13.2 Future works
Our method explicitly decomposes a complex scene into static and transient fields. However, transient elements are ephemeral and short-lived in the scenes. As a result, the absence of multi-view consistency eliminates the need to consider depth. Thus, deforming the 2D transient field of HybridGS [14], which consists of Gaussians with nine parameters, improves efficiency and reduces memory usage. Specifically, while we demonstrate the effectiveness of our methods on the Photo Tourism [27] and NeRF On-the-Go [22] datasets, another promising direction is extending our approach to large-scale dynamic scenes such as KITTI-360 [13] and NuScenes [1]. We leave this as future work.
Figure 18: Additional qualitative results from novel-view synthesis on the Photo Tourism dataset.
Figure 19: Additional qualitative results from novel-view synthesis on the NeRF On-the-go dataset.
Method Brandenburg Gate Sacre Coeur Trevi Fountain
PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓
3D-GS [7] 19.33 0.884 0.132 17.70 0.845 0.176 17.08 0.714 0.241
NeRF-W [17] 24.17 0.891 0.167 19.20 0.808 0.192 18.97 0.698 0.265
Ha-NeRF [18] 24.04 0.887 0.139 20.02 0.801 0.171 20.18 0.691 0.223
DeSplat [33] 25.04 0.92 0.142 20.14 0.868 0.178 23.31 0.775 0.226
CR-NeRF [40] 26.53 0.900 0.106 22.07 0.823 0.152 21.48 0.712 0.207
RobustNeRF [23] 25.79 0.923 0.094 20.94 0.852 0.137 23.58 0.785 0.170
IE-NeRF [32] 25.33 0.898 0.158 20.37 0.861 0.169 20.76 0.719 0.217
SWAG [3] 26.33 0.929 0.139 21.16 0.86 0.185 23.10 0.815 0.208
WildGaussian [10] 27.77 0.927 0.133 22.56 0.859 0.177 23.63 0.766 0.228
NexusSplats [28] 27.76 0.922 0.141 23.13 0.859 0.174 23.96 0.766 0.240
GS-W [44] 27.96 0.932 0.086 23.24 0.863 0.130 22.91 0.801 0.156
Wild-GS [37] 29.65 0.933 0.095 24.99 0.878 0.127 24.45 0.808 0.162
ForestSplats (Ours) 28.13 0.935 0.118 23.84 0.876 0.123 24.11 0.802 0.213
Table 9: Quantitative results on Photo Tourism dataset. The bold and underlined numbers indicate the best and second-best results.
Method Low Occlusion Medium Occlusion High Occlusion
Mountain Fountain Corner Patio Spot Patio-High
PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓
RobustNeRF [23] 17.54 0.496 0.383 15.65 0.318 0.576 23.04 0.764 0.244 20.39 0.718 0.251 20.65 0.625 0.391 20.54 0.578 0.366
Gaussian Opacity Field [43] 19.86 0.649 0.200 20.19 0.672 0.189 21.15 0.728 0.230 18.31 0.639 0.328 20.18 0.689 0.338 18.31 0.639 0.328
3D-GS [7] 19.40 0.638 0.213 19.96 0.659 0.185 20.90 0.713 0.241 17.48 0.704 0.199 20.77 0.693 0.316 17.29 0.604 0.363
Mip-Splatting [42] 20.70 0.661 0.169 20.37 0.662 0.187 21.53 0.739 0.241 15.58 0.491 0.536 20.03 0.683 0.324 15.58 0.491 0.536
GS-W [44] 19.43 0.596 0.299 20.06 0.723 0.274 22.17 0.793 0.155 19.90 0.681 0.260 17.13 0.608 0.409 19.90 0.681 0.260
NeRF On-the-go [22] 20.46 0.661 0.186 20.79 0.661 0.195 23.74 0.806 0.127 20.88 0.754 0.133 22.80 0.800 0.132 21.57 0.706 0.205
WildGaussians [10] 20.43 0.653 0.255 20.81 0.662 0.215 24.16 0.822 0.045 21.44 0.800 0.138 23.82 0.816 0.138 22.23 0.725 0.206
SpotLessSplats [24] 19.84 0.58 0.294 20.19 0.612 0.258 24.03 0.795 0.258 21.55 0.838 0.065 23.52 0.756 0.185 20.31 0.664 0.259
DeSplat [33] 19.59 0.710 0.170 20.27 0.680 0.170 26.05 0.880 0.090 20.89 0.810 0.110 26.07 0.900 0.090 22.59 0.840 0.120
HybridGS [14] 21.73 0.693 0.284 21.11 0.674 0.252 25.03 0.847 0.151 21.98 0.812 0.169 24.33 0.794 0.196 21.77 0.741 0.211
ForestSplats (Ours) 22.17 0.736 0.235 21.62 0.741 0.198 26.17 0.891 0.136 21.76 0.849 0.134 25.94 0.846 0.109 22.74 0.784 0.129
Table 10: Quantitative results on NeRF On-the-go dataset. The bold and underlined numbers indicate the best and second-best results. * indicates that the results are reproduced from the official code.
Figure 20: Visualization of appearance interpolation. We gradually interpolate from the source appearance to the target appearance.
Figure 21: Comparison of transient mask quality with existing methods on the Photo Tourism dataset.
Figure 22: Additional novel-view synthesis results for static and transient elements on the Photo Tourism and NeRF On-the-go datasets.
Figure 23: Comparison of image matching results on the NeRF On-the-go dataset with and without UAD across all scenes.
Figure 24: More qualitative results for each component in ForestSplats. DTF, SAM, MST, and UAD denote Deformable MLP, Superpixel-aware mask, Multi-stage training scheme, and Uncertainty-aware densification.
Figure 25: Our user study questionnaire. Each participant was shown an upper figure, which is a rendering video of several scenes using different methods. After watching the video, they were asked to answer the question below.
Notation Shape Definition
$\mathcal{\mu}$ $\mathbb{R}^{3}$ Position of Gausssians in 3DGS [7]
$\Sigma$ $\mathbb{R}^{3\times 3}$ Covariance matrix of Gaussians
$\alpha$ $\mathbb{R}^{1}$ Opacity of Gaussians
$c$ $\mathbb{R}^{3}$ View-dependent color of Gaussians
$W$ $\mathbb{R}^{3\times 3}$ Viewing transformation matrix
$J$ $\mathbb{R}^{2\times 3}$ Jacobian of the affine approximation of the projective transformation
$K$ Number of Gaussians contributed to the rendering
$C$ $\mathbb{R}^{3}$ The obtained pixel value after rendering
$M$ $\mathbb{R}^{1}$ Transient element mask
$\hat{I}$ $\mathbb{R}^{3\times H\times W}$ Rendered image
$I_{gt}$ $\mathbb{R}^{3\times H\times W}$ Ground truth image
$\hat{I}_{d}$ $\mathbb{R}^{3\times H\times W}$ Rendered image with transient elements
$\hat{I}_{s}$ $\mathbb{R}^{3\times H\times W}$ Rendered image with static elements
$t_{i}$ $\mathbb{R}^{32}$ Transient embedding per image $i$
$\Delta\mu$ $\mathbb{R}^{3}$ Offset for the Gaussian’s position
$\Delta s$ $\mathbb{R}^{3}$ Offset for the Gaussian’s scaling
$\Delta r$ $\mathbb{R}^{4}$ Offset for the Gaussian’s rotation
[…]
$\mathcal{M}_{O}$ Transient element mask from Unet $S_{\theta}$
$S_{\theta}$ Learnable CNN to mask transient elements
$\mathcal{S}$ The set where $I_{gt}$ is partitioned into superpixels
$s$ The elements of $\mathcal{S}$
$\mathcal{M}_{S}$ Superpixel-aware mask
$l^{(u)}$ $\mathbb{R}^{n}$ Uncertainty parameter for static Gaussians
$\mathbf{g}$ View-space positional gradient
$\mathcal{L}_{\mathrm{GS}}$ The loss item used in 3DGS [7]
$\mathcal{L}_{BCE}$ The loss to mitigate the undesired blending of Gaussians between static and transient fields
$\mathcal{L}_{\mathrm{init}}$ The loss to train the static field $\mathcal{G}_{s}$
$\mathcal{L}_{mid}$ The loss to train the Unet $S_{\theta}$ and transient field $\mathcal{G}_{d}$
$\lambda$ A weighting factor used in $\mathcal{L}_{GS}$
$f_{d}$ MLP predicting the attributes of transient Gaussians
$f_{c}$ MLP mapping the color of transient Gaussians
$\mathcal{G}_{s}$ Static field
$\mathcal{G}_{d}$ Deformable tranisent field
Table 11: For clarity and to avoid confusion in notation, we provided notation table.

### v3

@@ 6 Evaluation
As shown in Fig. 4 and Tab. 2, our method shows favorable performance compared to prior methods. In Fig. 4, we observe that CR-NeRF [38], 3D-GS [6], and GS-W [42] frequently fail to capture fine details in scenarios with appearance variations and struggle to handle transient elements. Unlike existing methods, WildGaussians [9] and Wild-GS [35] leverage semantic features from VFM [37, 18] to handle the transient elements. In contrast, our ForestSplats effectively decomposes transient elements from 2D scenes by leveraging transient field and a superpixel-aware mask without VFM, as shown in Tab. 2. More details are in the supplements.
@@
Comparison on the NeRF On-the-go Dataset. Furthermore, we also show the results for the NeRF On-the-go dataset [20]. Our method achieves competitive performance against existing SOTA methods across various occlusion levels without VFM in Tab. 1. Compared to RobustNeRF [21], and NeRF On-the-go [20], our method exhibits faster training and rendering speeds while mitigating occluders. RobustNeRF [21] considers photometric errors and the small size of the sampled patches to overlook complex structures and high occlusion scenarios. Specifically, WildGaussians [9] and SpotLessSplats [22] effectively remove transient elements by considering semantic features. Although leveraging semantic features facilitates handling transient elements, distinguishing semantic features similar to static elements remains challenging. However, our method explicitly decomposes static and transient elements across different levels of occluder scenarios, as shown in Fig. 5. Furthermore, compared to the prior methods without VFM, our method outperforms performance in terms of PSNR and SSIM.
@@
7 Ablation Study
Efficiency of Deformable Transient Field.
As shown in Tab. 3, the deformable transient field significantly reduces the memory usage (MB). Specifically in the Sacre Coeur, although transient embedding per image is considered, the deformable transient field achieves over 80% memory efficiency (MB). Compared to existing methods [31, 12], the deformable transient field adjusts the transient Gaussians based on per-scene optimization, showing efficient rendering for transient elements. Furthermore, we also analyze the performance based on the number of transient Gaussians as shown in Fig. 7-(a). We observe that the number of transient Gaussians primarily aims to occupy the regions with transient elements, which results in a slight effect on the representation of the static field. Moreover, we compare the transient field with and without the deformation MLP, as shown in Tab. 4. These results indicate that the deformable transient field not only significantly reduces memory usage (MB) but also facilitates scalability, particularly in large-scale scenarios.
@@ 7 Ablation Study
Figure 12: Ablation of uncertainty-aware densification (UAD) on opacity distribution and depth consistency. (a) Opacity distribution of Gaussians. (b) Consistency of depth. We visualize the effect of uncertainty-aware densification (UAD) on opacity and depth.
Effectiveness of superpixel-aware mask. We conducted an ablation study to assess the impact of a superpixel-aware mask in handling transient elements. As shown in Fig. 8, our method explicitly decomposes distractors from the static scene and represents transient objects in various scenarios, handling varying levels of occlusion. Furthermore, we analyze the effect of the number of superpixels in Fig. 9. According to the results, a small number of superpixels has minimal effect on the transient mask. Meanwhile, a large number of superpixels may lead to excessive masking of irrelevant areas. Thus, we set the number of segments to 20 as an optimal number. Moreover, we compare the quality of the transient mask in Fig. 10-(a). Existing methods [38, 42] rely on photometric errors. In contrast, superpixel-aware masks leverage superpixels, which consider both color and spatial information, to enhance mask quality without additional memory. Furthermore, the superpixel-aware mask serves as a plug-and-play module that combines with prior methods, showing improvement, as shown in Fig. 10-(b). In addition, we demonstrate a multi-stage training scheme that consistently improves performance in Fig. 7-(b).
Discussion on uncertainty-aware densification. We use uncertainty-aware densification (UAD) only in the static field to avoid the generation of static Gaussians at the boundaries of distractors during optimization. To analyze this, we examine an ablation study with and without UAD to see its effect on novel-view synthesis. As shown in Fig. 11, we observe that the centroids of static Gaussians are primarily positioned on static elements, mitigating the artifacts in novel view synthesis. In addition to, We analyze the effect of UAD on opacity and depth in Fig. 12. Our results show that UAD yields Gaussians exhibiting higher opacity values and improved depth consistency.
Analysis of each component. For further analysis, we evaluate each component of our method to demonstrate the effectiveness: Deformable transient field (DTF), Superpixel-aware mask (SAM), Multi-stage training scheme (MST), and Uncertainty-aware densification (UAD) as shown in Tab. 5. The results indicate that a superpixel-aware mask, which considers the photometric errors and the superpixels, combined with a multi-stage training scheme, consistently improves performance by effectively removing transient elements in scenes. Furthermore, we show qualitative results on the Brandenburg Gate and Corner scenes in Fig. 13.
@@ 8 Limitation and Conclusion
Although the transient field does not exhibit high quality due to deformable Gaussians, as shown in Fig. 8, the deformable transient field effectively captures transient elements while showing memory efficiency. Another limitation of our forestsplats is its difficulty in generalizing the number of superpixels due to the varying levels of occlusion in datasets, as depicted in Fig. 9.
@@
Conclusion. In this paper, we introduce ForestSplats, a novel framework that leverages a deformable transient field and a superpixel-aware mask to effectively remove distractors from static scenes. Furthermore, we propose a simple yet effective uncertainty-aware densification to enhance rendering quality and avoid generating static Gaussians on the boundaries of distractors during optimization. Extensive experiments on several datasets demonstrate that our method shows state-of-the-art performance compared to existing methods without a pre-trained model and shows significant memory efficiency to represent transient elements.